OpenStudy (gigi1103):

b. Write an expression equivalent to 4x + 12 using the distributive property.

OpenStudy (gigi1103):

@texaschic101

OpenStudy (phi):

can you write 12 as 4*3 ? yes so do that 4x+4*3

OpenStudy (gigi1103):

huh?

OpenStudy (phi):

now "factor out" 4 from each term (this is the opposite of "distributing" ) for example if you had 2(x+1) you would do 2*x + 2*1 or 2x+2 to undo that we write 2x+2 as 2*x + 2*1 and then "take out" the 2 that is in both terms. that leaves (x+1) and we put the 2 out front: 2(x+1)

OpenStudy (phi):

for your problem 4x+12 we first write it as 4*x + 4*3 now "take out" the 4 and put it "out front"

OpenStudy (gigi1103):

like 4(x *3)

OpenStudy (phi):

almost. it is x+3 inside the parens

OpenStudy (gigi1103):

so 4(x+3) ?

OpenStudy (phi):

yes. as a check we can "distribute the 4" 4(x+3) = 4*x + 4*3 = 4x +12 which shows 4(x+3) means the same thing as 4x+12 we started with 4x+12 and "went backwards" and got 4(x+3)

OpenStudy (gigi1103):

so is that my answer?

OpenStudy (phi):

yes

OpenStudy (gigi1103):

thanks!!